Business and finance. The demand equation for a certain type of printer is given by D= -200p + 35,000
The supply equation is predicted to be
S= -p2 + 400p - 20,000
Find the equilibrium price.

thanks

Answer by stanbon(75887)

Equilibrium means Demand = Supply:
----------------------
D= -200p + 35,000
The supply equation is predicted to be
S= -p2 + 400p - 20,000
-------------------------
-200p + 35,000 = -p2 + 400p - 20,000
p^2-600p+55000=0
p = [600+-sqrt140000)]/2
p = [600+-374.17]/2
p = 487.08 or p=112.915
==============
Cheers,
Stan H.